# Causal Climate Emulation with Bayesian Filtering

### v1

@@ Abstract
Traditional models of climate change use complex systems of coupled equations to simulate physical processes across the Earth system. These simulations are highly computationally expensive, limiting our predictions of climate change and analyses of its causes and effects. Machine learning has the potential to quickly emulate data from climate models, but current approaches are not able to incorporate physics-informed causal relationships. Here, we develop an interpretable climate model emulator based on causal representation learning. We derive a physics-informed approach including a Bayesian filter for stable long-term autoregressive emulation. We demonstrate that our emulator learns accurate climate dynamics, and we show the importance of each one of its components on a realistic synthetic dataset and data from two widely deployed climate models.
@@ 1 Introduction
In order to respond to climate change, it is necessary to understand future climates as precisely as possible, under different scenarios of anthropogenic greenhouse gas emissions. Earth system models (ESMs), which couple together complex numerical models simulating different components of the Earth system, are the primary tool for making these projections, but require solving vast numbers of equations representing physical processes and are thus computationally expensive (Balaji et al., 2017; Mansfield et al., 2020). This severely limits the number of future scenarios that can be simulated and attribution studies investigating the influence of climate change on extreme events. Reduced physical models, such as Simple Climate Models ((Leach et al., 2021), are commonly used to speed up the process, but at the expense of oversimplifying necessary complex processes.
@@
  • We develop a novel machine learning algorithm for causal emulation of climate model dynamics, incorporating (i) physics-informed losses that ensure invariant properties of the data are preserved, (ii) a Bayesian filter that allows for stable long-term emulation when autoregressively rolling out predictions.
@@ 2.1 Causal learning for time series
Causal representation learning aims to address this gap. Under certain conditions, it is possible to disentangle relevant latent variables from high-dimensional observed data (Hyvarinen and Morioka, 2017; Hyvarinen et al., 2019; Khemakhem et al., 2020). Recent work demonstrated that a latent representation and a causal graph between latent variables can be learned simultaneously (Lachapelle et al., 2020; Schölkopf et al., 2021) via a fully differentiable framework (Zheng et al., 2018; Brouillard et al., 2020). CDSD (Brouillard et al., 2024) aims to recover both the latent variables and the temporal causal graph over these latents from a high-dimensional time series, where the causal graph is shown to be identifiable under the strong single-parent decoding assumption – i.e. each observed variable is mapped to a single latent, while each latent can be the parent of multiple observed variables. This assumption is well-suited to modeling long-range climate dynamics and teleconnections, where variations in distinct regions around the world can causally affect the climate in other regions.
@@ 3 Methods
To meet the challenges of learning causal structure, physical realism, and interpretability, we introduce the PICABU model (Physics-Informed Causal Approach with Bayesian Uncertainty). We describe each aspect of this approach here, and the overall pipeline is illustrated in Figure 1, showing the latent representation, the loss used during training, the Bayesian filter, and counterfactual experiments.
Figure 1: High-level schematic of the PICABU pipeline. The key features of the PICABU pipeline are illustrated: the latent embeddings under single-parent assumption, the learned causal graph over these latents, the loss used during training, the Bayesian filter allowing for stable autoregressive rollouts, and the possibility of counterfactual experiments.
@@ 3.1 Core framework
Our framework leverages CDSD (Brouillard et al., 2024), which jointly learns latent variables $\mathbf{Z}$ and a mapping from $\mathbf{Z}$ to the high-dimensional observations $\mathbf{X}$, and proves identifiability of the causal graph under the single-parent assumption (for details, see Appendix B). This latent-to-observation graph is parameterized by a matrix $W$ and neural networks. Latents at time $t$ are mapped to latents at previous timesteps through a directed acyclic graph (DAG) and neural networks. The parameters are learned by maximizing an evidence lower bound (ELBO). The augmented Lagrangian method (ALM, Nocedal and Wright, 2006) is used to optimize the loss under the constraint $\mathcal{C}_{\text{single-parent}}^{\{\lambda,\mu\}}$ and ensure that $W$ is orthogonal and satisfy the assumptions needed for identifiability. A penalty $\mathcal{P}_{\text{sparsity}}^{\lambda}$ is added to the loss to sparsify the DAG and retain only causally-relevant links between latents.
@@
3.2 Physics-informed loss functions
To improve the predictions of the learned model for longer rollouts, we include a second loss term, which is a measure of probabilistic next timestep prediction accuracy, the continuous ranked probability score (CRPS, Matheson and Winkler (1976); Hersbach (2000)).
@@ 3.2 Physics-informed loss functions
$\mathbf{X}$ is the true next timestep from the training data and $F(\hat{\mathbf{X}})$ is the cumulative distribution function of the predicted next timestep. Furthermore, to ensure that the model preserves the spatiotemporal structure of the data, we incorporate loss terms measuring the $\ell^{1}$ difference between the true and inferred spatial spectrum of the data, and between the true and inferred temporal spectra for each grid cell:
@@
We find that the coefficients of the penalty terms and the initial values of the ALM for the constraints need to be carefully chosen to ensure effective training of the model. Appendix N discusses the coefficients and hyperparameter choices, and gives the values used in our experiments.
@@ 3.3 Bayesian filtering for stable autoregressive rollouts
Once trained, our model gives a distribution over latent variables given latents at previous timesteps: $p(\mathbf{z}^{t}|\mathbf{z}^{<t})$ (learned transition model). To perform long-term climate prediction, we use an autoregressive model that predicts the next timestep sequentially. Rather than predicting at each step $\mathbf{z}^{t}$ from the mean of $p(\mathbf{z}^{<t})$, we perform recursive Bayesian estimation and sample $N$ values from $p(\mathbf{z}^{t}|\mathbf{z}^{<t})$, to preserve the full distribution through time. Furthermore, as each latent is mapped to a prediction in the observation space, we add a particle filtering step. For each sample $\mathbf{z}^{\leq t}_{n}$ at time $\leq t$, we sample $R$ values from $p(\mathbf{z}^{t+1}|\mathbf{z}_{n}^{\leq t})$, and assign a likelihood to each of them, which encode whether the samples are mapped to an observation satisfying the desired statistics of the data. By keeping the sample with the highest likelihood value, we prevent error accumulation or convergence to the mean and obtain a set of $N$ stable trajectories from our autoregressive rollout.
When performing recursive Bayesian estimation for a state-space model, we typically aim to compute the posterior distribution for the latents given a sequence of observations:
@@
To sample from this posterior distribution, we can estimate $p(\mathbf{z}^{t}|\mathbf{x}^{<t})$ (learned encoder model), but unfortunately we do not observe $\mathbf{x}^{t}$ as we are performing prediction and cannot sample from $p(\mathbf{x}^{t}|\mathbf{z}^{t})$. We propose using the Fourier spatial spectrum at each timestep as a proxy for our observations: we assume that it is constant through time and treat its value as the true observation. We then compute $p(\tilde{\mathbf{x}}|\mathbf{z}^{t})$ where $\tilde{\mathbf{x}}$ is the constant spatial spectrum of the climate variable over Earth, obtained from available data. We then sample from the posterior distribution $p(\mathbf{z}^{t}|\mathbf{\tilde{x}^{t}},\mathbf{z}^{<t})$, and eliminate samples with low importance score: if a sample corresponds to a spatial spectrum that is very far from the true spatial spectrum, it will be assigned a low probability and thus will be discarded. We represent $p(\mathbf{\tilde{x}}|\mathbf{z}^{t})$ using a Laplace distribution $\mathcal{L}(\mathbf{\tilde{x}}^{t},\tilde{\sigma})$. The mean $\mathbf{\tilde{x}}^{t}$ is obtained by decoding $\mathbf{z}^{t}$ and mapping them to observations $\mathbf{x}^{t}$, which are then Fourier-transformed to obtain $\mathbf{\tilde{x}}^{t}$. The variance $\tilde{\sigma}$ is estimated directly from the observations and assumed constant through time.
One could, in practice, use any known statistic of the data. We focus on the spatial spectrum as predicting an accurate spatial spectrum is necessary to prevent models from propagating errors during rollouts (Chattopadhyay et al., 2024; Guan et al., 2024). This approach could be applied to many statistics known to be constant or invariant through time, including conserved quantities (White et al., 2024). Once the spatial spectrum, assumed fixed, is computed by taking the mean of spatial spectra across observed NorESM2 data, our autoregressive rollout works as given in Algorithm 1.
@@
  Input: Observations $\mathbf{x}^{\leq T}$, trained encoder $p(\mathbf{z}^{\leq t}|\mathbf{x}^{\leq t})$, trained decoder $\mathbf{x}=f(\mathbf{z})$, learned transition model $p(\mathbf{z}^{t}|\mathbf{z}^{<t})$, ground-truth spatial spectrum $\mathbf{\tilde{x}}$ with standard deviation $\tilde{\sigma}$, number of sampled trajectories $N$ and sample size $R$, prediction time range $m$
  Initialization: Get $N$ samples $\mathbf{z}_{n}^{\leq T}\sim p(\mathbf{z}^{\leq T}|\mathbf{x}^{\leq T})$
@@
     For each $n\leq N,r\leq R$, decode $\mathbf{x}_{n,r}^{T+i}=f(\mathbf{z}_{n,r}^{T+i})$ and use FFT to get the associated spectrum $\tilde{x}_{n,r}^{T+i}$ and weights $w_{n,r}^{T+i}=\mathcal{L}(\mathbf{\tilde{x}}|\mathbf{\tilde{x}}_{n,r}^{T+i},\tilde{\sigma})$
@@
  Output: $N$ latent trajectories $\{\mathbf{z}_{n\leq N}^{T\leq t\leq T+m}\}$
@@ 4.1 Recovering a known causal graph
The spatially averaged vector autoregressive (SAVAR) model (Tibau et al., 2022) allows the creation of time series with known underlying causal relations, designed to benchmark causal discovery methods. The model generates data following an autoregressive process, with modes of variability interacting through defined connections mimicking those in the climate system (Nowack et al., 2020). It exhibits two essential properties of climate models: spatial aggregation (each grid cell is influenced by neighboring cells) and both local and long-range dependencies. Although it is a simplified model that does not fully capture real-world climate systems, it provides a controlled setting with a known ground-truth causal structure. Thus, SAVAR allows us to rigorously assess the ability of PICABU to accurately recover causal relationships from high-dimensional data.
Data. We generate several datasets with $N=4,25,\text{or }100$ latents, and four varying levels of difficulty. In the “easy” dataset, latents only depend on themselves at one previous timestep. In the “med-easy” dataset, they additionally depend on each latent at a previous timestep with probability $p=1/(N-1)$. This probability grows to $p=2/(N-1)$ for the “med-hard” and $p=1/2$ for the “hard” dataset. The causal dependencies are restricted to the five previous timesteps. The expected number of edges in the true graph is thus $M=N+N\cdot(N-1)\cdot p$. The strength values of the autoregressive causal links are drawn from a beta distribution $\mathcal{B}(4,8)$ with mean $1/3$ and standard deviation $0.13$, and normalized. The generative process of the SAVAR data is linear. Noise is added to the datasets with variance equal to the signal variance. SAVAR datasets are illustrated in Appendix C.
Models. We compare PICABU to two baseline methods: CDSD (Brouillard et al., 2024) and Varimax-PCMCI (V-PCMCI) (Runge et al., 2015). We report in Table 1 the F1-score $\frac{2}{\text{recall}^{-1}+\text{precision}^{-1}}$ for all three methods. The F1-score ranges from 0 to 1, indicating how well the true graph is learned, with 1 indicating a perfect graph. For all methods, we set the number of latent variables equal to the true latent dimension of the SAVAR datasets and use linear transition and decoding models. We run PICABU without the spectral components of the loss (i.e. $\lambda_{\text{s}}=\lambda_{\text{t}}=0$) as the data is not complex enough to require additional physics-informed penalties. For a fair comparison, all models retain the true number of edges $M$ in their graphs. As V-PCMCI scales poorly, we do not report results on the 100 latent datasets due to the extensive computational resources this experiment would require. The details of the training procedure and hyperparameter search are given in Section C.1.
@@
Results. PICABU achieves high F1-scores on all datasets of various dimensionalities and difficulties, and performs slightly better than CDSD. Both methods disentangle the latents very well in all datasets. The physics-informed loss is not needed here to achieve high accuracy, as the SAVAR datasets are of relatively low complexity. When tested on climate model data, CDSD predictions diverge (see §4.2), showing the fundamental importance of the physics-informed loss terms when modeling real climate model data. Another benefit of PICABU is the possibility to control the final sparsity of the graph directly via a constraint. CDSD, in contrast, requires careful early-stopping or a thorough hyperparameter search over its sparsity penalty coefficient.
V-PCMCI achieves lower accuracy, even on datasets of lower complexity or latent dimensionality. We argue that because the latents are learned using Varimax-PCA before the causal connections are learned, it might not disentangle the latents correctly, especially as noise increases (Section C.2).
@@ 4.2 Pre-industrial climate model emulation
Data. To assess performance in emulating widely used climate models, we evaluate PICABU on monthly NorESM2 (Seland et al., 2020) and CESM2-FV2 (Danabasoglu et al., 2020) surface temperature data at a resolution of 250 km, re-gridded to an icosahedral grid. In this section, we focus on pre-industrial data, using 800 years of data provided by two NorESM2 simulations initialized with different initial conditions. We normalize and deseasonalize the data by subtracting the monthly mean and dividing by the monthly standard deviation at each grid point. Figure 8 shows an example of six successive timesteps. The 800 years of data are split into 90% train and 10% test sets. Results for NorESM2 are reported below; for CESM2-FV2, refer to Appendix I.
Models. PICABU is trained to take five consecutive months as input and predict the next month. An example prediction is shown in Figure 2. V-PCMCI learns a causal graph and a mapping from latents to observables, but not the full transition model. We enhance it by fitting, for each latent, a regression on the previous latents it causally depends on, according to the learned PCMCI graph. This allows us to predict the next timestep, given the five previous timesteps. We also train a Vision Transformer (Dosovitskiy et al., 2021) with a positional encoding adapted from Nguyen et al. (2023), referred to as “ViT + pos. encoding”. Details on this model are given in Appendix D. Furthermore, we perform a study in which we individually ablate each term in PICABU’s loss. After training, we sample initial conditions from the test set and, for each initial condition, autoregressively project 100 years forward. PICABU, ablated models, and V-PCMCI use our Bayesian filter, whereas “ViT + pos. encoding” does not, as it is deterministic.
@@
Results. As climate model data is chaotic, high-dimensional, and driven by complex spatiotemporal dynamics, we seek a model that has both low mean bias over long rollouts and captures the spatiotemporal variability of the data. To capture these properties, we evaluate the mean, standard deviation, range, and log-spectral distance (LSD) of key climate variables, with the goal being for ML models to come as close as possible to the ground truth values for each of these statistics. We compute these values for two climate variables of great relevance to long-term climate dynamics: the normalized global mean surface temperature (GMST) and the Niño3.4 index, which represents the El Niño Southern Oscillation (ENSO) (the most important natural source of variability in the climate system). The GMST and Niño3.4 index statistics are computed over ten 50-year periods from NorESM2 pre-industrial data. In Table 2, we compare values for the ground-truth NorESM2 data and the models PICABU, V-PCMCI, “ViT + pos. encoding”, and ablations.
Overall, we find that PICABU comes closest to matching the ground truth across all statistics of the GMST and Niño3.4 index. By contrast, CDSD diverges and is unable to fit the data at all, V-PCMCI gives predictions with unrealistically low spatiotemporal variability (as captured in the std. dev. and range being far from ground-truth values) on both climate variables, and the ViT with positional encoding exhibits overly high variability, especially for GMST. Our tests also affirm the value of the different PICABU components. Removing the CRPS term leads to a higher range in GMST, while removing the spectral term leads to much higher bias. Removing the orthogonality constraint leads to poorer performance for GMST and comes at the expense of identifiability and interpretability. PICABU without the Bayesian filter or sparsity term both diverge.
@@
Table 2: PICABU shows a small bias and captures climate variability better than comparable methods. The left section shows the mean, standard deviation, range, and log spectral distance (LSD) of the GMST, and the same for the Niño3.4 index on the right. The first row shows the values of the metrics for the ground-truth NorESM2, followed by comparison methods, and then PICABU ablations. For all metrics, being closer to the ground truth is better.
We also consider the power spectral density (PSD) Froyland et al. (2021) of the Niño3.4 index and GMST for NorESM2 data and the different models (Figure 3). We find that PICABU captures the temporal dynamics of Niño3.4 index and GMST better than all other models. In particular, the power spectrum of the Niño3.4 index shows a peak at a period of three years, and the power spectrum of PICABU closely approximates the ground truth around this peak, illustrating that the model has learned accurate climate dynamics, while the other models do not perform as well (Figure 9).
Figure 3: PICABU learns accurate temporal variability for ENSO, and outperforms other methods in learning GMST variability. We run PICABU for ten different 50-year emulations, and compute the mean and standard deviation of the spectra, doing the same for ten different 50-year periods of NorESM2 data. On the left, the power spectra for the Niño3.4 index, for PICABU, the ViT, V-PCMCI, and ground-truth data (NorESM2) are shown. The same is shown for GMST on the right.
However, we note that, while outperforming the other models, PICABU exhibits too much power at very high frequencies for Niño3.4 and GMST (Figure 11), suggesting room for further improvement in emulating large-scale variability at short timescales. Similar plots for the ablated PICABU models are given in Appendix F (Figure 10), with all ablations showing a decrease in performance as expected.
Results for two further important known patterns of variability, the Indian Ocean Dipole (IOD) and the Atlantic Multidecadal Oscillation (AMO) (Saji et al., 1999; Kerr, 2000; McCarthy et al., 2015), are shown in Appendix H, and are consistent with the above results for ENSO and GMST. Additionally, performance on the CESM2-FV2 climate model (Appendix I) is similar to the results for NorESM2 data. To complement Table 2, the average annual temperature range for each grid cell is calculated for both NorESM2 and for PICABU, and compared in Figure 13. The spatial pattern of the annual range of projected temperatures is reasonable, but has a greater range in some regions (Appendix G).

### v2

@@ Abstract
Traditional models of climate change use complex systems of coupled equations to simulate physical processes across the Earth system. These simulations are highly computationally expensive, limiting our predictions of climate change and analyses of its causes and effects. Machine learning has the potential to quickly emulate data from climate models, but current approaches are not able to incorporate physically-based causal relationships. Here, we develop an interpretable climate model emulator based on causal representation learning. We derive a novel approach including a Bayesian filter for stable long-term autoregressive emulation. We demonstrate that our emulator learns accurate climate dynamics, and we show the importance of each one of its components on a realistic synthetic dataset and data from two widely deployed climate models.
@@
now at the European Centre for Medium-Range Weather Forecasts, UK
@@ 1 Introduction
In order to respond to climate change, it is necessary to understand future climates as precisely as possible, under different scenarios of anthropogenic greenhouse gas emissions. Earth system models (ESMs), which couple together complex numerical models simulating different components of the Earth system, are the primary tool for making these projections, but require solving vast numbers of equations representing physical processes and are thus computationally expensive (Balaji et al., 2017; Mansfield et al., 2020). This severely limits the number of future scenarios that can be simulated and attribution studies investigating the influence of climate change on extreme events. Reduced physical models, such as Simple Climate Models (Leach et al., 2021), are commonly used to speed up the process, but at the expense of oversimplifying necessary complex processes.
@@
  • We develop a novel machine learning algorithm for causal emulation of climate model dynamics, incorporating (i) additional losses that ensure invariant properties of the data are preserved, (ii) a Bayesian filter that allows for stable long-term emulation when autoregressively rolling out predictions.
@@ 2.1 Causal learning for time series
Causal representation learning aims to address this gap. Under certain conditions, it is possible to disentangle relevant latent variables from high-dimensional observed data (Hyvarinen and Morioka, 2017; Hyvarinen et al., 2019; Khemakhem et al., 2020). Recent work demonstrated that a latent representation and a causal graph between latent variables can be learned simultaneously (Lachapelle et al., 2020; Schölkopf et al., 2021) via a fully differentiable framework (Zheng et al., 2018; Brouillard et al., 2020). CDSD (Brouillard et al., 2024; Boussard et al., 2023) aims to recover both the latent variables and the temporal causal graph over these latents from a high-dimensional time series, where the causal graph is shown to be identifiable under the strong single-parent decoding assumption – i.e. each observed variable is mapped to a single latent, while each latent can be the parent of multiple observed variables.
In our setting, this assumption imposes that the latents correspond to single climate modes in a geographic region which interact through the causal graph. While strong, the assumption is well-suited to modeling long-range climate dynamics and teleconnections, where variations in distinct regions around the world can causally affect the climate in other regions. The learned latents mirror simplified climate indices (e.g., scalar indices representing climate variables in fixed regions) that climate scientists widely rely on to describe large-scale atmospheric dynamics. Climate scientists often use simple representations of complex processes for interpretable, parsimonious description of large-scale phenomena. Our work parallels such methods closely. Furthermore, the single-parent assumption is a principled first step towards interpretable and trustworthy data-driven climate models. It makes our tool interpretable for climate scientists and gives theoretical guarantees of causal identifiability. This allows exploration of the effect of interventions through counterfactual experiments, which we hope will be useful for attribution studies by domain scientists.
@@ 3 Methods
To meet the challenges of learning causal structure, physical realism, and interpretability, we introduce the PICABU model (Physics-Informed Causal Approach with Bayesian Uncertainty). We describe each aspect of this approach here, and the overall pipeline is illustrated in Figure˜1, showing the latent representation, the loss used during training, the Bayesian filter, and counterfactual experiments.
Figure 1: High-level schematic of the PICABU pipeline. The key features of the PICABU pipeline are illustrated: the latent embeddings under the single-parent assumption, the learned causal graph over these latents, the loss used during training, the Bayesian filter allowing for stable autoregressive rollouts, and the possibility of counterfactual experiments.
@@ 3.1 Core framework
Our framework leverages CDSD (Brouillard et al., 2024), which jointly learns latent variables $\mathbf{Z}$ and a mapping from $\mathbf{Z}$ to the high-dimensional observations $\mathbf{X}$, and proves identifiability of the causal graph under the single-parent assumption (for details, see Appendix˜B). This latent-to-observation graph is parameterized by a matrix $W$ and neural networks. Latents at time $t$ are mapped to latents at previous timesteps through a directed acyclic graph (DAG) and neural networks. The parameters are learned by maximizing an evidence lower bound (ELBO). The augmented Lagrangian method (ALM, Nocedal and Wright, 2006) is used to optimize the loss under the constraint $\mathcal{C}_{\text{single-parent}}^{\{\lambda,\mu\}}$ and ensure that $W$ is orthogonal and satisfies the assumptions needed for identifiability. A penalty $\mathcal{P}_{\text{sparsity}}^{\lambda}$ is added to the loss to sparsify the DAG and retain only causally-relevant links between latents.
@@
3.2 Additional loss functions
To improve the predictions of the learned model for longer rollouts, we include a second loss term, the continuous ranked probability score (CRPS; Matheson and Winkler (1976); Hersbach (2000)), a measure of probabilistic next-timestep prediction accuracy.
$\mathcal{L}_{\text{CRPS}}=\int_{-\infty}^{\infty}\left[F(\hat{\mathbf{X}})-\mathbf{1}_{\{\mathbf{X}\leq\hat{\mathbf{X}}\}}\right]^{2}d\hat{\mathbf{X}}$ (2)
@@ 3.2 Additional loss functions
We find that the coefficients of the penalty terms and the initial values of the ALM for the constraints need to be carefully chosen to ensure effective model training. Appendix˜O discusses the coefficients and hyperparameter choices, and gives the values used in our experiments.
@@ 3.3 Bayesian filtering for stable autoregressive rollouts
Once trained, our model gives a distribution over latent variables given the latents at previous timesteps: $p(\mathbf{z}^{t}|\mathbf{z}^{<t})$ (the learned transition model). To perform long-term climate prediction, we use an autoregressive model that predicts the next timestep sequentially. Rather than predicting at each step $\mathbf{z}^{t}$ from the mean of $p(\mathbf{z}^{<t})$, we perform recursive Bayesian estimation and sample $N$ values from $p(\mathbf{z}^{t}|\mathbf{z}^{<t})$, to preserve the full distribution through time. Furthermore, as each latent is mapped to a prediction in the observation space, we add a particle filtering step. For each sample $\mathbf{z}^{\leq t}_{n}$ at time $\leq t$, we sample $R$ values from $p(\mathbf{z}^{t+1}|\mathbf{z}_{n}^{\leq t})$, and assign a likelihood to each of them, which encode whether the samples are mapped to an observation satisfying the desired statistics of the data. By keeping the sample with the highest likelihood value, we prevent error accumulation or convergence to the mean and obtain a set of $N$ stable trajectories from our autoregressive rollout.
When performing recursive Bayesian estimation for a state-space model, we typically aim to compute the posterior distribution for the latents given a sequence of observations as follows:
@@
[…]
In practice, one could use any known statistic of the data. We focus on the spatial spectrum as predicting an accurate spatial spectrum is necessary to prevent models from propagating errors during rollouts (Chattopadhyay et al., 2024; Guan et al., 2024). This approach could be applied to many statistics known to be constant or invariant through time, including conserved quantities (White et al., 2024). Once the spatial spectrum, assumed fixed, is computed by taking the mean of spatial spectra across observed NorESM2 data, our autoregressive rollout proceeds as given in Algorithm˜1.
@@
Input: Observations $\mathbf{x}^{\leq T}$, trained encoder $p(\mathbf{z}^{\leq t}|\mathbf{x}^{\leq t})$, trained decoder $\mathbf{x}=f(\mathbf{z})$, learned transition model $p(\mathbf{z}^{t}|\mathbf{z}^{<t})$, ground truth spatial spectrum $\mathbf{\tilde{x}}$ with standard deviation $\tilde{\sigma}$, number of sampled trajectories NN and sample size $R$, prediction time range $m$
Initialization: Get NN samples $\mathbf{z}_{n}^{\leq T}\sim p(\mathbf{z}^{\leq T}|\mathbf{x}^{\leq T})$
@@
[…]
@@
Output: NN latent trajectories $\{\mathbf{z}_{n\leq N}^{T\leq t\leq T+m}\}$
@@ 4.1 Recovering a known causal graph
The spatially averaged vector autoregressive (SAVAR) model (Tibau et al., 2022) allows the creation of time series with known underlying causal relations, designed to benchmark causal discovery methods. The model generates data following an autoregressive process, with modes of variability interacting through defined connections mimicking those in the climate system (Nowack et al., 2020). It exhibits two essential properties of climate models: spatial aggregation (each grid cell is influenced by neighboring cells) and both local and long-range dependencies. Although it is a simplified model that does not fully capture real-world climate systems, it provides a controlled setting with a known ground truth causal structure. Thus, SAVAR allows us to rigorously assess the ability of PICABU to accurately recover causal relationships from high-dimensional data.
Data. We generate several datasets with $N=4,25,\text{or }100$ latents, and four varying levels of difficulty. In the “easy” dataset, latents depend only on themselves at one previous timestep. In the “med-easy” dataset, they additionally depend on each latent at a previous timestep with probability p=1/(N1)p=1/(N-1). This probability grows to p=2/(N1)p=2/(N-1) for the “med-hard” and $p=1/2$ for the “hard” dataset. The causal dependencies are restricted to the five previous timesteps. The expected number of edges in the true graph is thus $M=N+N\cdot(N-1)\cdot p$. The strength values of the autoregressive causal links are drawn from a beta distribution $\mathcal{B}(4,8)$ with mean $1/3$ and standard deviation $0.13$, chosen to keep most link strengths moderate and away from 0 or 1, and are then normalized. The generative process of the SAVAR data is linear. Noise is added to the datasets with variance equal to the signal variance. SAVAR datasets are illustrated in Appendix˜C.
Models. We compare PICABU to two baseline methods: CDSD (Brouillard et al., 2024) and Varimax-PCMCI (V-PCMCI) (Runge et al., 2015). We report in Table˜1 the F1-score $\frac{2}{\text{recall}^{-1}+\text{precision}^{-1}}$ for all three methods. The F1-score ranges from 0 to 1, indicating how well the true graph is learned, with 1 indicating a perfect graph. For all methods, we set the number of latent variables equal to the true latent dimension of the SAVAR datasets and use linear transition and decoding models. We run PICABU without the spectral components of the loss (i.e. $\lambda_{\text{s}}=\lambda_{\text{t}}=0$) as the data is not complex enough to require additional penalties. For a fair comparison, all models retain the true number of edges $M$ in their graphs. As V-PCMCI scales poorly, we do not report results on the 100 latent datasets due to the extensive computational resources this experiment would require. The details of the training procedure and hyperparameter search are given in Section˜C.2.
@@
Results. PICABU achieves high F1-scores on all datasets of various dimensionalities and difficulties, and performs slightly better than CDSD. Both methods disentangle the latents very well in all datasets. The additional losses are not needed here to achieve high accuracy, as the SAVAR datasets are of relatively low complexity. When tested on climate model data, CDSD predictions diverge (see §4.2), showing the fundamental importance of the additional loss terms when modeling real climate model data. Another benefit of PICABU is the possibility to control the final sparsity of the graph directly via a constraint. CDSD, in contrast, requires careful early-stopping or a thorough hyperparameter search over its sparsity penalty coefficient.
V-PCMCI achieves lower accuracy, even on datasets of lower complexity or latent dimensionality. We argue that because the latents are learned using Varimax-PCA before the causal connections are learned, it might not disentangle the latents correctly, especially as noise increases (Section˜C.3).
@@ 4.2 Pre-industrial climate model emulation
Data. To assess performance in emulating widely used climate models, we evaluate PICABU on monthly NorESM2 (Seland et al., 2020) and CESM2-FV2 (Danabasoglu et al., 2020) surface temperature data at a resolution of 250 km, re-gridded to an icosahedral grid. In this section, we focus on pre-industrial data, using 800 years of data provided by two NorESM2 simulations initialized with different initial conditions. We normalize and deseasonalize the data by subtracting the monthly mean and dividing by the monthly standard deviation at each grid point. Figure˜10 shows an example of six successive timesteps. The 800 years of data are split into 90% train and 10% test sets. Results for NorESM2 are reported below; for CESM2-FV2, refer to Appendix˜I.
Models. PICABU is trained to take five consecutive months as input and predict the next month. An example prediction is shown in Figure˜2. V-PCMCI learns a causal graph and a mapping from latents to observables, but not the full transition model. We enhance it by fitting, for each latent, a regression on the previous latents it causally depends on, according to the learned PCMCI graph. This allows us to predict the next timestep, given the five previous timesteps. We also train a Vision Transformer (Dosovitskiy et al., 2021) with a positional encoding adapted from Nguyen et al. (2023), referred to as “ViT + pos. encoding”. Details on this model are given in Appendix˜D. Furthermore, we perform a study in which we individually ablate each term in PICABU’s loss. After training, we sample initial conditions from the test set and, for each initial condition, autoregressively project 100 years forward. PICABU, ablated models, and V-PCMCI use our Bayesian filter, whereas “ViT + pos. encoding” does not, as it is deterministic.
@@
Results. As climate model data is chaotic, high-dimensional, and driven by complex spatiotemporal dynamics, we seek a model that has both low mean bias over long rollouts and captures the spatiotemporal variability of the data. To capture these properties, we evaluate the mean, standard deviation, range, and log-spectral distance (LSD) of key climate variables, with the goal being for ML models to come as close as possible to the ground truth values for each of these statistics. We compute these values for two climate variables of great relevance to long-term climate dynamics: the normalized global mean surface temperature (GMST) and the Niño3.4 index, which represents the El Niño Southern Oscillation (ENSO) (the most important natural source of variability in the climate system). The GMST and Niño3.4 index statistics are computed over ten 50-year periods from NorESM2 pre-industrial data. In Table˜2, we compare values for the ground truth NorESM2 data and the models PICABU, V-PCMCI, “ViT + pos. encoding”, and ablations.
Overall, we find that PICABU comes closest to matching the ground truth across all statistics for the GMST and Niño3.4 indices. By contrast, CDSD diverges and is unable to fit the data at all. V-PCMCI gives predictions with unrealistically low spatiotemporal variability (as captured in the std. dev. and range being far from ground truth values) for both climate variables, and the ViT with positional encoding exhibits overly high variability, especially for GMST. Our tests also affirm the value of the different PICABU components. Removing the CRPS term leads to a higher range in GMST, while removing the spectral term leads to much higher bias. Removing the orthogonality constraint leads to poorer performance for GMST and comes at the expense of identifiability and interpretability. PICABU without the Bayesian filter or sparsity term both diverge.
@@
Table 2: PICABU shows a small bias and captures climate variability better than comparable methods. The left section shows the mean, standard deviation, range, and log spectral distance (LSD) of the GMST, and the same for the Niño3.4 index on the right. The first row shows the values of the metrics for the ground truth NorESM2, followed by comparison methods, and then PICABU ablations. For all metrics, being closer to the ground truth is better.
We also consider the power spectral density (PSD, Froyland et al. (2021)) of the Niño3.4 index and GMST for NorESM2 data and the different models (Figure˜3). We find that PICABU captures the temporal dynamics of Niño3.4 index and GMST better than all other models. In particular, the power spectrum of the Niño3.4 index shows a peak at a period of three years, and the power spectrum of PICABU closely approximates the ground truth around this peak, illustrating that the model has learned accurate climate dynamics, while the other models do not perform as well (Figure˜11).
Figure 3: PICABU learns accurate temporal variability for ENSO, and outperforms other methods in learning GMST variability. We run PICABU for ten different 50-year emulations, and compute the mean and standard deviation of the spectra, doing the same for ten different 50-year periods of NorESM2 data. On the left, the power spectra for the Niño3.4 index, for PICABU, the ViT, V-PCMCI, and ground truth data (NorESM2) are shown. The same is shown for GMST on the right.
Figure 4: PICABU outperforms ablated models in learning ENSO and GMST variability. We run PICABU for ten different 50-year emulations, and compute the mean and standard deviation of the spectra, doing the same for ten different 50-year periods of NorESM2 data. On the left, the power spectra for the Niño3.4 index, for PICABU and ablations, and ground truth data (NorESM2) are shown. On the right, the same is shown for GMST.
However, we note that, while outperforming the other models, PICABU exhibits too much power at very high frequencies for Niño3.4 and GMST (Appendix˜F, Figure˜12), suggesting room for further improvement in emulating large-scale variability at short timescales. Similar plots for the ablated PICABU models are given in Figure˜4, with all ablations showing a decrease in performance, illustrating the importance of the additional loss terms.
Results for two further important known patterns of variability, the Indian Ocean Dipole (IOD) and the Atlantic Multidecadal Oscillation (AMO) (Saji et al., 1999; Kerr, 2000; McCarthy et al., 2015), are shown in Appendix˜H, and are consistent with the above results for ENSO and GMST. Additionally, performance on the CESM2-FV2 climate model (Appendix˜I) is similar to the results for NorESM2 data. To complement Table˜2, the average annual temperature range for each grid cell is calculated for both NorESM2 and for PICABU, and compared in Figure˜14. The spatial pattern of the annual range of projected temperatures is reasonable, but has a greater range in some regions (Appendix˜G).